eviant:

PLS HELP!!!! The midpoint of KL is M(–8, 1). One endpoint is K(–6, 5). Find the coordinates of the other endpoint L.

eviant:

@Shadow @bm717

bm717:

1 attachment
Shadow:

\[M = ( \frac{ x _{1} + x _{2}}{ 2 } , \frac{ y _{1} + y _{2} }{ 2 })\]

bm717:

Frick hold on x'D

bm717:

LMAO STOP

bm717:

BYE

Shadow:

Just a formula (:

bm717:

Die.

eviant:

so, the midpoint formula right?

Shadow:

Yes, just input your points, add, then divide.

eviant:

so, -10, -3?

Shadow:

Well actually this is one endpoint and a midpoint, so this is a bit different. Lol

Shadow:

What did you do?

eviant:

I did -8-6=14 to the second power, which is 3.7

eviant:

wait, I'm using the wrong formula, sorry

eviant:

8-6=14, 14/2=7

eviant:

then, 1+5=6, 6/2=3

eviant:

so, it should be 7, 3 right?

Shadow:

You're skipping the negatives, and that's now how you would approach it.

Shadow:

We need to find the other endpoint. The formula I posted below is when you have both endpoints, you need to restructure it a bit for this problem.

eviant:

what do you mean?

Shadow:

The midpoint of KL is M(–8, 1). One endpoint is K(–6, 5). Find the coordinates of the other endpoint L. This is asking for the "other endpoint L" \[M = ( \frac{ x _{1} + x _{2}}{ 2 } , \frac{ y _{1} + y _{2} }{ 2 })\] This solves for the midpoint when you have both of the endpoints. We already have the midpoint, we need the other endpoint.

Shadow:

But where's what you can do. \[-8 = \frac{ x _{1} + x _{2}}{ 2 } , 1 = \frac{ y _{1} + y _{2} }{ 2 }\]

Shadow:

Input the values for the endpoint you have, and solve for the x and y value for the other end point.

Shadow:

Let me know if you need help with the steps.